The minimum and maximum distance of a point(1,2)from the ellipse 4x^2+9y^2+8x-36y+4=0are L and G .Then the value of G-L is-

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

(x-1)2+(y-2)2,

4(x+1)2+3(y-2)2=12

(x-1)2+((12-4(x+1)2)/3)=f(x)

find maximum and minimum of f(x)